If F(x) 1/39(3x-2)2;0<_x_<3


O: elsewhere


1. Verify that F(x) is a PDF


2. find E(x) and Var(x)


1
Expert's answer
2022-06-06T02:46:55-0400

1.


"\displaystyle\int_{-\infin}^{\infin}F(x)dx=\displaystyle\int_{0}^{3}\dfrac{(3x-2)^2}{39}dx"

"=[\dfrac{(3x-2)^3}{3(3)(39)}]\begin{matrix}
 3\\
 0
\end{matrix}=\dfrac{343+8}{351}=1, True"

2.


"E(X)=\displaystyle\int_{-\infin}^{\infin}F(x)xdx=\displaystyle\int_{0}^{3}\dfrac{x(3x-2)^2}{39}dx"

"=\displaystyle\int_{0}^{3}\dfrac{9x^3-12x^2+4x}{39}dx"

"=[\dfrac{1}{39}(\dfrac{9x^4}{4}-4x^3+2x^2)]\begin{matrix}
 3\\
 0
\end{matrix}"

"=\dfrac{1}{39}(\dfrac{729}{4}-108+18-0)"

"=\dfrac{123}{52}"



"E(X^2)=\displaystyle\int_{-\infin}^{\infin}F(x)x^2dx=\displaystyle\int_{0}^{3}\dfrac{x^2(3x-2)^2}{39}dx"

"=\displaystyle\int_{0}^{3}\dfrac{9x^4-12x^3+4x^2}{39}dx"

"=[\dfrac{1}{39}(\dfrac{9x^5}{5}-3x^4+\dfrac{4x^3}{3})]\begin{matrix}
 3\\
 0
\end{matrix}"

"=\dfrac{1}{39}(\dfrac{2187}{5}-243+36-0)"

"=\dfrac{384}{65}"

"Var(X)=E(X^2)-(E(X))^2"

"=\dfrac{384}{65}-(\dfrac{123}{52})^2=\dfrac{4227}{13520}"